A rectangular field is to be enclosed on four sides with a fence. Fencing costs $4 per foot for two opposite sides, and $7 per foot for the other two sides. Find the dimensions of the field of area 740 ft 2 that would be the cheapest to enclose.




A. 36 ft @ $4 by 20.6 ft @ $7




B. 20.6 ft @ $4 by 36 ft @ $7




C. 47.6 ft @ $4 by 15.5 ft @ $7




D. 15.5 ft @ $4 by 47.6 ft @ $7

1
Expert's answer
2021-11-01T19:04:28-0400



Length=X

Width=y

Area of rectangle="740ft^2"

"XY=740.....(1)"

Length of rectangle is 4per foot and width is 7per foot

"C(x,y)=4x+7y\\=4x+7x(\frac{740}{X})\\=4x+\frac{5180}{x}"

"xy=740\\y=\frac{740}{x}"

For the cheapest value

"c(x)=0\\\frac{d}{dx}(4x+\frac{5180}{x})=0\\
\frac{5180}{x^2}=4"

"x^2=\frac{5180}{4}\\
X=36feet"

"xy=740\\y=\frac{740}{35.906}=20.6feet"


Answer is


A. 36 ft @ $4 by 20.6 ft @ $7